# Written by AAAI Press Staff1

### v2

@@ Abstract
Abstract
1 Introduction and Related Work
@@
Graph partitioning is a fundamental problem in network analysis with numerous real-world applications in various domains such as system design in online social networks (pujol2009divide), dynamic ride-sharing in transportation systems (tafreshian2020trip), VLSI design (kahng2011vlsi), and preventing cascading failure in power grids (li2005strategic). The goal of graph partitioning is to divide a given graph into disjoint subsets where nodes within each subset exhibit strong internal connections while having limited connections with nodes in other subsets. Generally speaking, the aim is to find somewhat balanced partitions while minimizing the number of edges across partitions.
@@
Several graph partitioning formulations have been studied in the literature, mostly in the form of discrete optimization (karypis1997multilevel; karypis1998multilevelk; karypis1999multilevel; andersen2006local; chung2007four). The majority of the formulations are NP-hard and thus the proposed solutions are either heuristics or algorithms with approximate solutions (karypis1997multilevel). Recently, there have been attempts to solve graph partitioning problem via neural approaches. The neural approaches have a distinct advantage that they can utilize node features. Node features supply additional information and provide contextual insights that may improve the accuracy of graph partitioning. For instance, (wang2017mgae) has recognized the importance of incorporating node attributes in graph clustering where nodes are partitioned into disjoint groups. Note that there are existing neural approaches (khalil2017learning; gcomb; kool2018attention; joshi2019efficient) to solve other NP-hard graph combinatorial problems (e.g., minimum vertex cover). However these methods are not generic enough to solve graph partitioning.
@@
In this paper, we build a single framework to solve several graph partitioning problems. One of the most relevant to our work is the method DMoN by muller2023graph. This method designs a neural architecture for cluster assignments and use a modularity-based objective function for optimizing these assignments. Another method, that is relevant to our work is GAP, which is an unsupervised learning method to solve the balanced graph partitioning problem (nazi2019gap). It proposes a differentiable loss function for partitioning based on a continuous relaxation of the normalized cut formulation. Deep-MinCut being an unsupervised approach learns both node embeddings and the community structures simultaneously where the objective is to minimize the mincut loss (duong2023deep). Another method solves the multicut problem where the number of partitions is not an input to the problem (jung2022learning). The idea is to construct a reformulation of the multicut ILP constraints to a polynomial program as a loss function. However, the problem formulation is different than the generic normalized cut or mincut problem. Finally,  (gatti2022graph) solves the normalized cut problem only for the case where the number of partitions is exactly two. Nevertheless, these neural approaches for the graph partitioning problem often do not use node features and only limited to a distinct partitioning objective. Here, we point out notable drawbacks that we address in our framework.
@@
  • Non-inductivity to partition count: The number of partitions required to segment a graph is an input parameter. Hence, it is important for a learned model to generalize to any partition count without retraining. Existing neural approaches are non-inductive to the number of partitions, i.e they can only infer on number of partitions on which they are trained. Additionally, it is worth noting that the optimal number of partitions is often unknown beforehand. Hence, it is a common practice to experiment with different partition counts and evaluate their impact on the partitioning objective. While the existing methods (muller2023graph; nazi2019gap; gatti2022graph) assume that the number of partitions ($k$) is known beforehand, our proposed method can generalize to any $k$.
@@
  • Non-generalizability to different cut functions: Multiple objective functions for graph cut have been studied in the partitioning literature. The optimal objective function hinges upon the subsequent application in question. For instance, the two most relevant studies, DMoN (muller2023graph) and GAP (nazi2019gap) focus on maximizing modularity and minimizing normalized cut respectively. Our framework is generic and can solve different partitioning objectives.
@@
  • Assumption of differential objective function: Existing neural approaches assume the objective function to be differentiable. As we illustrate in § 2, the assumption does not always hold in the real-world. For instance, the sparsest cut (chawla2006hardness) and balanced cut (nazi2019gap) objectives are not differentiable.
1.1 Contributions:
@@
  • Empirical Assessment: We perform comprehensive experiments employing real-world datasets, evaluating NeuroCUT across four distinct graph partitioning objectives. Our empirical investigation substantiates the efficacy of NeuroCUT in partitioning tasks, showcasing its robustness across a spectrum of objective functions. We also demonstrate the capacity of NeuroCUT to generalize effectively to partition sizes that it has not encountered during training.
2 Problem Formulation
@@
Definition 2 (Cut)
A cut $\mathcal{C}=(\mathcal{S},\mathcal{T})$ is a partition of $\mathcal{V}$ into two subsets $\mathcal{S}$ and $\mathcal{T}$. The cut-set of $\mathcal{C}=(\mathcal{S},\mathcal{T})$ is the set $(u,v)\in\mathcal{E}|u\in\mathcal{S},v\in\mathcal{T}$ of edges that have one endpoint in $\mathcal{S}$ and the other endpoint in $\mathcal{T}$.
Definition 3 (Graph Partitioning)
@@
  1. 1.
    $k$-MinCut (saran1995finding): Partition a graph into $k$ partitions such that the total number of edges across partitions is minimized.
@@
  2. 2.
    Normalized Cut (shi2000normalized): The $k$-MinCut criteria favors cutting small sets of isolated nodes in the graph. To avoid this unnatural bias for partitioning out small sets of points, normalised cut computes the cut cost as a fraction of the total edge connections to all the nodes in the graph.
    $\operatorname{Ncut}(\mathcal{\mathcal{P}}):=\sum_{l=1}^{|\mathcal{P}|}\frac{|\operatorname{cut}\left(P_{l},\overline{P_{l}}\right)|}{\operatorname{vol}\left(P_{l},\mathcal{V}\right)}$ (2)
@@
  3. 3.
    Balanced Cuts (nazi2019gap): Balanced cut favours partitions of equal sizes so an extra term that indicates the squared distance from equal sized partition is added to normalised cuts.
    $\operatorname{Balanced-Cuts}(\mathcal{\mathcal{P}}):=\sum_{l=1}^{|\mathcal{P}|}\frac{|\operatorname{cut}\left(P_{l},\overline{P_{l}}\right)|}{\operatorname{vol}\left(P_{l},\mathcal{V}\right)}+\frac{(|P_{l}|-|\mathcal{V}|/k)^{2}}{|\mathcal{V}|^{2}}$ (3)
@@
  4. 4.
    Sparsest Cuts (chawla2006hardness): Two-way sparsest cuts minimize the cut edges relative to the number of nodes in the smaller partition. We generalize it to $k$-way sparsest cuts by summing up the value for all the partitions. The intuition behind sparsest cuts is that any partition should neither be very large nor very small.
    $\phi(S,\bar{S})$ $=\frac{\operatorname{cut}(S,\bar{S})}{\min(|S|,|\bar{S}|)}$ (4)
    $\operatorname{Sparsest-Cuts}(\mathcal{\mathcal{P}})$ $=\sum_{l=1}^{|\mathcal{P}|}\phi(P_{l},\overline{P_{l}})$ (5)
Problem 1 (Learning to Partition Graph )
@@
Figure 1: Architecture of NeuroCUT. First, the initial partitioning of the graph is performed based on node raw features and positional embeddings. These embeddings are refined using GNN to infuse toplogical information from neighborhood. At each step a node is selected and its partitioned is updated. During the training phase the GNN parameters are updated and hence embeddings are re-computed. During inference, the GNN is called only once to compute the embeddings of the nodes of the graph.
@@
$\mathcal{G}$
Graph
$\mathcal{V}$
Node set
$e$
Edge $e\in\mathcal{E}$
$\mathcal{E}$
Edge set
X
Feature matrix containing raw node features
$\mathcal{N}_{v}$
Neighboring nodes of node $v$
$Obj(\mathcal{G},\mathcal{P})$
Objective function based upon graph $\mathcal{G}$ and its partitioning $\mathcal{P}$
$k$
Number of partitions
$\mathcal{P}^{t}$
Partitioning at time $t$
${P}^{t}_{i}$
$i^{th}$ partition at time $t$
$\overline{{P}_{i}}$
Set of nodes that are not in the $i^{th}$ partition
$\mathcal{S}^{t}$
State of system at step $t$
$\operatorname{PART}(\mathcal{P}^{t},v)$
Partition of node $v$ at time $t$
$\mathcal{S}^{t}$
State representation of Partitions and Graph at step $t$
$\mathbf{pos}(v)$
Positional embedding of node $v$
$\mathbf{emb_{init}}(v)$
Initial embedding of node $v$
$\alpha$
Number of anchor nodes for lipschitz embedding
$T$
Length of trajectory
$\pi$
Policy function
Table 1: Notations used in the paper
3 NeuroCUT: Proposed Methodology
Fig. 1 describes the framework of NeuroCUT. For a given input graph $\mathcal{G}$, we first construct the initial partitions of nodes using a clustering based approach. Subsequently, a message-passing Gnn embeds the nodes of the graph ensuring inductivity to different graph sizes. Next, the assignment of nodes to partitions proceeds in a two-phased strategy. We first select a node to change its partition and then we choose a suitable partition for the selected node. Further, to ensure inductivity on the number of partitions, we decouple the number of partitions from the direct output representations of the model. This decoupling allows us to query the model to an unseen number of partitions. After a node’s partition is updated, the reward with respect to change in partitioning objective value is computed and the parameters of the policy are optimized. The reward is learned through reinforcement learning (RL) (sutton2018reinforcement).
@@ 3 NeuroCUT: Proposed Methodology
3.1 State: Initialization & Positional Encoding
Initialization. Instead of directly starting from empty partitions, we perform a warm start operation that clusters the nodes of the graph to obtain the initial graph partitions. The graph is clustered into $k$ clusters based on their raw features and positional embeddings (discussed below). We use K-means (macqueen1967some) algorithm for this task. The details of the clustering are present in Appendix A.2.
Positional Encoding (Embeddings). Given that the partitioning objectives are NP-hard mainly due to the combinatorial nature of the graph structure, we look for representations that capture the location of a node in the graph. Positional encodings provide an idea of the position in space of a given node within the graph. Two nodes that are closer in the graph, should be closer in the embedding space. Towards this, we use Lipschitz Embedding (bourgain1985lipschitz). Let $\mathcal{A}=\{a_{1},\cdots,a_{\alpha}\}\subseteq\mathcal{V}$ be a randomly selected subset of $\alpha$ nodes. We call them anchor nodes. From each anchor node $i$, the walker starts a random walk (bianchini2005inside) and jumps to a neighboring node $j$ with a transition probability ($\mathbf{W}_{ij}$) governed by the transition probability matrix $\mathbf{W}\in\mathbb{R}^{|V|\times|V|}$. Furthermore, at each step, with probability $c$ the walker jumps to a neighboring node $j$ and returns to the node $i$ with $1-c$. Let $\vec{r}_{ij}$ corresponds to the probability of the random walker starting from node $i$ and reaching node $j$.
$\vec{r}_{i}=c\tilde{\mathbf{W}}\vec{r}_{i}+(1-c)\vec{e}_{i}$ (6)
@@ 3.1 State: Initialization & Positional Encoding
$\mathbf{pos}(u)=[r_{1u},r_{2u},\cdots,r_{\alpha u}]$ (7)
@@
$\mathbf{emb_{init}}(u)=\textbf{X}[u]\;\;\mathbin{\|}\;\;\mathbf{pos}(u)$ (8)
@@
$S^{t}=\{S_{1}^{t},S_{2}^{t},\cdots S_{k}^{t}:\;S_{i}^{t}{=}\{\mathbf{emb_{init}}(v)\;\;\forall v\in P_{i}^{t}\}\}$ (9)
@@
3.2 Action: Selection of Nodes and Partitions

### v3

# NeuroCUT: A Neural Approach for Robust Graph Partitioning

@@ Abstract.
Graph Partitioning, Min Cut, Robustness, Versatile Partitioning Objectives, Inductive learning, GNN, Reinforcement learning.
copyright: acmcopyrightjournalyear: 2023doi: 10.1145/3637528.3671815journalyear: 2024copyright: acmlicensedconference: 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining ; August 25–29, 2024; Barcelona, Spainbooktitle: 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, 2024, Barcelona, Spainisbn: 979-8-4007-0490-1/24/08doi: 10.1145/3637528.3671815ccs: Computing methodologies Machine learningccs: Computing methodologies Artificial intelligenceccs: Mathematics of computing Graph algorithmsccs: Mathematics of computing Graph algorithms
1. Introduction
@@
Graph partitioning is a fundamental problem in network analysis with numerous real-world applications in various domains such as system design in online social networks (Pujol et al., 2009), dynamic ride-sharing in transportation systems (Tafreshian and Masoud, 2020), VLSI design (Kahng et al., 2011), and preventing cascading failure in power grids (Li et al., 2005). The goal of graph partitioning is to divide a given graph into disjoint subsets where nodes within each subset exhibit strong internal connections while having limited connections with nodes in other subsets. Generally speaking, the aim is to find somewhat balanced partitions while minimizing the number of edges across partitions.
1.1. Related Work
@@
Several graph partitioning formulations have been studied in the literature, mostly in the form of discrete optimization (Karypis et al., 1997; Karypis and Kumar, 1998, 1999; Andersen et al., 2006; Chung, 2007). The majority of the formulations are NP-hard and thus the proposed solutions are either heuristics or algorithms with approximate solutions (Karypis et al., 1997). Among these, two widely used methods are Spectral Clustering (Ng et al., 2001) and hMETIS (Karypis et al., 1997). Spectral Clustering partitions a graph into clusters based on the eigenvectors of a similarity matrix derived from the graph. hMETIS is a hypergraph partitioning method that divides a graph into clusters by maximizing intra-cluster similarity while minimizing inter-cluster similarity. However, such techniques are confined to specific objective functions and are unable to leverage the available node features in the graph. Recently, there have been attempts to solve graph partitioning problem via neural approaches. The neural approaches have a distinct advantage that they can utilize node features. Node features supply additional information and provide contextual insights that may improve the accuracy of graph partitioning. For instance, (Wang et al., 2017) has recognized the importance of incorporating node attributes in graph clustering where nodes are partitioned into disjoint groups. Note that there are existing neural approaches (Khalil et al., 2017; Manchanda et al., 2020; Kool et al., 2018; Manchanda et al., 2022, 2024; Joshi et al., 2019; Tian et al., 2024; Gasse et al., 2019; Ranjan et al., 2022; Manchanda and Ranu, 2023; Goyal et al., 2020) to solve other NP-hard graph combinatorial problems (e.g., minimum vertex cover). However these methods are not generic enough to solve graph partitioning. For example, although S2VDQN (Khalil et al., 2017) learns to solve the Maxcut problem, it is designed for the specific case of bi-partitioning. This hinders its applicability to the target problem setup of $k$-way partitioning [^1] Additional related work is present in Appendix LABEL:app:additional_rel_work.
@@
In this paper, we build a single framework to solve several graph partitioning problems. One of the most relevant to our work is the method DMoN by Tsitsulin et al. (2023a). This method designs a neural architecture for cluster assignments and use a modularity-based objective function for optimizing these assignments. Another method, that is relevant to our work is GAP, which is an unsupervised learning method to solve the balanced graph partitioning problem (Nazi et al., 2019). It proposes a differentiable loss function for partitioning based on a continuous relaxation of the normalized cut formulation. Deep-MinCut being an unsupervised approach learns both node embeddings and the community structures simultaneously where the objective is to minimize the mincut loss (Duong et al., 2023). Another method solves the multicut problem where the number of partitions is not an input to the problem (Jung and Keuper, 2022). The idea is to construct a reformulation of the multicut ILP constraints to a polynomial program as a loss function. However, the problem formulation is different than the generic normalized cut or mincut problem. Another related work is DGCluster (Bhowmick et al., 2024), which proposes to solve the attributed graph clustering problem while maximizing modularity when $k$ is not known beforehand. Finally,  (Gatti et al., 2022) solves the normalized cut problem only for the case where the number of partitions is exactly two. Nevertheless, these neural approaches for the graph partitioning problem often do not use node features and only limited to a distinct partitioning objective. Here, we point out notable drawbacks that we address in our framework.
@@
  • Non-inductivity to partition count: The number of partitions required to segment a graph is an input parameter. Hence, it is important for a learned model to generalize to any partition count without retraining. In $k$-way graph partitioning, a model demonstrates inductivity to the number of partitions when it can infer on varying partition numbers without specific training for each. Existing neural approaches are non-inductive to the number of partitions, i.e they can only infer on number of partitions on which they are trained. Additionally, it is worth noting that the optimal number of partitions is often unknown beforehand. This capability is crucial in practical applications like chip design, where graph partitioning optimizes logic cell placement by dividing netlists (circuits) into smaller subgraphs, aiding independent placement. As the optimal partition count is frequently unknown in advance, experimenting with different partition numbers is a common practice. Hence, it is a common practice to experiment with different partition counts and evaluate their impact on the partitioning objective. While the existing methods (Tsitsulin et al., 2023a; Nazi et al., 2019; Gatti et al., 2022) assume that the number of partitions ($k$) is known beforehand, our proposed method can generalize to any $k$.
@@
  • Non-generalizability to different cut functions: Multiple objective functions for graph cut have been studied in the partitioning literature. The optimal objective function hinges upon the subsequent application in question. For instance, the two most relevant studies, DMoN (Tsitsulin et al., 2023a) and GAP (Nazi et al., 2019) focus on maximizing modularity and minimizing normalized cut respectively. Our framework is generic and can solve different partitioning objectives.
@@
  • Assumption of differential objective function: Existing neural approaches assume the objective function to be differentiable. As we illustrate in § 2, the assumption does not always hold in the real-world. For instance, the sparsest cut (Chawla et al., 2006) and balanced cut (Nazi et al., 2019) objectives are not differentiable.
1.2. Our Contributions
@@
  • Empirical Assessment: We perform comprehensive experiments employing real-world datasets, evaluating NeuroCUT across four distinct graph partitioning objectives. Our empirical investigation substantiates the efficacy of NeuroCUT in partitioning tasks, showcasing its robustness across a spectrum of objective functions. We also demonstrate the capability of NeuroCUT to generalize effectively to partition sizes that have not been seen during training.
2. Problem Formulation
@@
Definition 2 (Cut).
A cut $\mathcal{C}=(\mathcal{S},\mathcal{T})$ is a partition of $\mathcal{V}$ into two subsets $\mathcal{S}$ and $\mathcal{T}$. The cut-set of $\mathcal{C}=(\mathcal{S},\mathcal{T})$ is the set ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\{}(u,v)\in\mathcal{E}|u\in\mathcal{S},v\in\mathcal{T}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\}}$ of edges that have one endpoint in $\mathcal{S}$ and the other endpoint in $\mathcal{T}$.
Definition 3 (Graph Partitioning).
@@
  1. (1)
    $k$-MinCut (Saran and Vazirani, 1995): Partition a graph into $k$ partitions such that the total number of edges across partitions is minimized.
@@
  2. (2)
    Normalized Cut (Shi and Malik, 2000): The $k$-MinCut criteria favors cutting small sets of isolated nodes in the graph. To avoid this unnatural bias for partitioning out small sets of points, normalised cut computes the cut cost as a fraction of the total edge connections to all the nodes in the graph.
    (2) $\operatorname{Ncut}(\mathcal{\mathcal{P}}):=\sum_{l=1}^{|\mathcal{P}|}\frac{|\operatorname{cut}\left(P_{l},\overline{P_{l}}\right)|}{\operatorname{vol}\left(P_{l},\mathcal{V}\right)}$
@@
  3. (3)
    Balanced Cut (Nazi et al., 2019): Balanced cut favours partitions of equal sizes so an extra term that indicates the squared distance from equal sized partition is added to normalized cuts.
    (3) $\operatorname{Balanced-Cuts}(\mathcal{\mathcal{P}}):=\sum_{l=1}^{|\mathcal{P}|}\frac{|\operatorname{cut}\left(P_{l},\overline{P_{l}}\right)|}{\operatorname{vol}\left(P_{l},\mathcal{V}\right)}+\frac{(|P_{l}|-|\mathcal{V}|/k)^{2}}{|\mathcal{V}|^{2}}$
@@
  4. (4)
    Sparsest Cut (Chawla et al., 2006): Two-way sparsest cuts minimize the cut edges relative to the number of nodes in the smaller partition. We generalize it to $k$-way sparsest cuts by summing up the value for all the partitions. The intuition behind sparsest cuts is that any partition should neither be very large nor very small.
    (4) $\operatorname{Sparsest-Cuts}(\mathcal{\mathcal{P}})$ $=\sum_{l=1}^{|\mathcal{P}|}\phi(P_{l},\overline{P_{l}})$
    (5) $where\;\;\phi(S,\bar{S})$ $=\frac{\operatorname{cut}(S,\bar{S})}{\min(|S|,|\bar{S}|)}$
Problem 1 (Learning to Partition Graph ).
@@
Figure 1. Architecture of NeuroCUT. First, the initial partitioning of the graph is performed based on node features and positional embeddings. These embeddings are refined using GNN to infuse toplogical information from neighborhood. At each step a node is selected and its partitioned is updated. During training the GNN parameters are updated and hence embeddings are re-computed. During inference, the GNN is called only once to compute the embeddings of the nodes of the graph.
Table 1. Notations used in the paper
@@
$\mathcal{G}$ Graph
$\mathcal{V}$ Node set
$e$ Edge $e\in\mathcal{E}$
$\mathcal{E}$ Edge set
X Feature matrix containing raw node features
$\mathcal{N}_{v}$ Neighboring nodes of node $v$
$Obj(\mathcal{G},\mathcal{P})$ Objective function based upon graph $\mathcal{G}$ and its partitioning $\mathcal{P}$
$k$ Number of partitions
$\mathcal{P}^{t}$ Partitioning at time $t$
${P}^{t}_{i}$ $i^{th}$ partition at time $t$
$\overline{{P}_{i}}$ Set of nodes that are not in the $i^{th}$ partition
$\mathcal{S}^{t}$ State of system at step $t$
$\operatorname{PART}(\mathcal{P}^{t},v)$ Partition of node $v$ at time $t$
$\mathcal{S}^{t}$ State representation of Partitions and Graph at step $t$
$\mathbf{pos}(v)$ Positional embedding of node $v$
$\mathbf{emb_{init}}(v)$ Initial embedding of node $v$
$\alpha$ Number of anchor nodes for lipschitz embedding
$T$ Length of trajectory
$\pi$ Policy function
3. NeuroCUT: Proposed Methodology
Fig. 1 describes the framework of NeuroCUT. For a given input graph $\mathcal{G}$, we first construct the initial partitions of nodes using a clustering based approach. Subsequently, a message-passing Gnn embeds the nodes of the graph ensuring inductivity to different graph sizes. Next, the assignment of nodes to partitions proceeds in a two-phased strategy. We first select a node to change its partition and then we choose a suitable partition for the selected node. Further, to ensure inductivity on the number of partitions, we decouple the number of partitions from the direct output representations of the model. This decoupling allows us to query the model to an unseen number of partitions. After a node’s partition is updated, the reward with respect to change in partitioning objective value is computed and the parameters of the policy are optimized. The reward is learned through reinforcement learning (RL) (Sutton and Barto, 2018).
@@ 3. NeuroCUT: Proposed Methodology
3.1. State: Initialization & Positional Encoding
Initialization. Instead of directly starting from empty partitions, we perform a warm start operation that clusters the nodes of the graph to obtain the initial graph partitions. The graph is clustered into $k$ clusters based on their raw features and positional embeddings (discussed below). We use K-means (MacQueen et al., 1967) algorithm for this task. The details of the clustering are present in Appendix A.1.
Positional Encoding (Embeddings). Given that the partitioning objectives are NP-hard mainly due to the combinatorial nature of the graph structure, we look for representations that capture the location of a node in the graph. Positional encodings provide an idea of the position in space of a given node within the graph. Two nodes that are closer in the graph, should be closer in the embedding space. Towards this, we use Lipschitz Embedding (Bourgain, 1985). Let $\mathcal{A}=\{a_{1},\cdots,a_{\alpha}\}\subseteq\mathcal{V}$ be a randomly selected subset of $\alpha$ nodes. We call them anchor nodes. From each anchor node $i$, the walker starts a random walk (Bianchini et al., 2005) and jumps to a neighboring node $j$ with a transition probability ($\mathbf{W}_{ij}$) governed by the transition probability matrix $\mathbf{W}\in\mathbb{R}^{|V|\times|V|}$. Furthermore, at each step, with probability $c$ the walker jumps to a neighboring node $j$ and returns to the node $i$ with $1-c$. Let $\vec{r}_{ij}$ corresponds to the probability of the random walker starting from node $i$ and reaching node $j$.
(6) $\vec{r}_{i}=c\tilde{\mathbf{W}}\vec{r}_{i}+(1-c)\vec{e}_{i}$
@@ 3.1. State: Initialization & Positional Encoding
(7) $\mathbf{pos}(u)=[r_{1u},r_{2u},\cdots,r_{\alpha u}]$
@@
(8) $\mathbf{emb_{init}}(u)=\textbf{X}[u]\;\;\mathbin{\|}\;\;\mathbf{pos}(u)$
@@
(9) $S^{t}=\{S_{1}^{t},S_{2}^{t},\cdots S_{k}^{t}:\;S_{i}^{t}{=}\{\mathbf{emb_{init}}(v)\;\;\forall v\in P_{i}^{t}\}\}$
@@
3.2. Action: Selection of Nodes and Partitions